Present Value of Bonds Payable; Discount Pinder Co. produces and sells high-quality video equipment. To finance its operations, Pinder Co. issued $25,000,000 of five-year, 7% bonds, with interest payable semiannually, at a market (effective) interest rate of 9%.

Determine the present value of the bonds payable, using the present value tables in Exhibit 8 and Exhibit 10.

Round to the nearest dollar. $
Answer #1
(note there will be slight difference in issue price calculation ,kindly use the factors as given
inn your question to get exact answer)
Bond characterstics Amount
1-a) Principal 25,000,000
interest            875,000
Market interest rate 4.5%
periods to maturity 10
issue price 23,021,880 answer
Calculation of bond issue price
Where
i= 4.50%
t= 10
principal * PV of $1 at 4.5% for 10 yrs =
25,000,000 * 0.64393        = 16098250
interest * PV of ordinary annuity at 4.5%=
875000 * 7.91272 = 6923630
bond issue price 23021880